OpenStudy (anonymous):

Find two consecutive odd integers whose sum is 72. Which equation would be used to solve this problem? (HINT: The equation has been simplified some.) 2x = 72 2x + 2 = 72 x + 2 = 72

OpenStudy (anonymous):

i would use common sense

OpenStudy (anonymous):

(2*x+1)+(2x+1+1)=72 i think hold on i check

OpenStudy (anonymous):

\(33+35=68\) too low

OpenStudy (anonymous):

try \(35+37\)

OpenStudy (anonymous):

the nimrod who wrote this problem wants you to say \( 2x + 2 = 72 \)

OpenStudy (anonymous):

@satellite73 @timo86m Thanks!

OpenStudy (anonymous):

subtract \(2\) and get \(2x=70\) so \(x=35\)

OpenStudy (anonymous):

this is really bad, treats people like morons there is no "one way" to solve this problem